John left work and traveled towards New York at average speed of 33 miles per hour Michael left two hours later and traveled in the same direction but with an average speed of 55 miles per hour how long did John travel until Micah caught up?
Answer by ankor@dixie-net.com(22740)

John left work and traveled towards New York at average speed of 33 miles per hour
Michael left two hours later and traveled in the same direction but with an average speed of 55 miles per hour
how long did John travel until Micah caught up?
:
Let t = travel time of J when M caught up with him
then
(t-2) = travel time of M when he catches J
:
When this happens, they will both have traveled the same distance
Write a dist equation; d = speed * time
55(t-2) = 33t
55t - 110 = 33t
55t - 33t = 110
22t = 110
t = 110/22
t = 5 hrs, J traveled when caught by M
:
:
We can confirm this solution, find the actual dist each traveled
Should be the same
55 * 3 = 165 Traveled 2 hrs less
33 * 5 = 165